OpenStudy (anonymous):

Express with radical signs. Please show me the steps! b^-1/2

OpenStudy (anonymous):

\[b^{\frac{ -1 }{ 2 }}=\frac{ 1 }{b ^{\frac{ 1 }{ 2 }}}=\frac{ 1 }{ \sqrt{b} }\]

OpenStudy (anonymous):

cool :)